OpenStudy (anonymous):

A line passes through point (3, 7) and has a slope of 3/4. The equation of the line is blank . If point A(x, 5) lies on the line, the value of x is blank .

OpenStudy (texaschic101):

y = mx + b slope(m) = 3/4 (3,7)...x = 3 and y = 7 now sub 7 = 3/4(3) + b 7 = 9/4 + b 7 - 9/4 = b 28/4 - 9/4 = b 19/4 = b so your equation is : y = 3/4x + 19/4 Now we have points (x,5)....where y = 5. So we sub 5 in for y in our equation to find x 5 = 3/4x + 19/4 5 - 19/4 = 3/4x 20/4 - 19/4 = 3/4x --- multiply by 4 to get rid of fractions 20 - 19 = 3x 1 = 3x 1/3 = x lets check.. y = 3/4x + 19/4 (1/3, 5)...x = 1/3 and y = 5 5 = 3/4(1/3) + 19/4 5 = 1/4 + 19/4 5 = 20/4 5 = 5 (correct) points are (1/3, 5)